OpenStudy (anonymous):

WHAT IS THE DIFFRENCE OF 1/2 & 4/5

OpenStudy (anonymous):

common denominators.....1/2=5/10....and 4/5=8/10.....so the difference would be 5/10 - 8/10 which = -3/10 or -.3